# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v2

@@ 7 $N_{\text{intra}}^{+}-N_{\text{inter}}^{+}$ 2 0 -2 0 -2
8 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}$1 1 -1 1 1
9 $-\,N_{\text{intra}}^{-}-N_{\text{inter}}^{+}$ ($-\,f^{\text{MinDisagree}}$) -1 -1 -3 -1 -2
10 $N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ ($-f^{\text{MinCut}}$) 0 0 -2 0 -1
@@ 13 $N_{\text{intra}}^{+}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ 2 1 -2 1 -1
14 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$1 0 -3 0 -1
@@
Now, consider a clustering with $k$ non-neutral clusters, where all intra-cluster similarities are $+1$ and all inter-cluster similarities are $-1$. If an unassigned object $i\in V$ has similarity $+1$ to all others, the max correlation objective (Eq. 11) assigns it to the largest non-neutral cluster, while the minimum cut objective (Eq. 12) keeps it neutral. For $k>2$, the full objective (Eq. 1) places $i$ in the largest non-neutral cluster if its size exceeds the sum of all others; otherwise, it remains neutral. Conversely, if object $i$ has similarity $-1$ to all others, max correlation keeps it neutral, whereas minimum cut assigns it to the smallest non-neutral cluster. For $k>2$, the full objective may assign $i$ as neutral or non-neutral depending on cluster sizes.
@@
For $k=2$, the full objective always increases (or remains constant) when an object is made non-neutral, aligning it with max agreement. To see this, consider a clustering with $k=2$ non-neutral clusters, and let $M_{i,m}=\sum_{j\in S_{m}}A_{i,j}$ be the total similarity of object $i$ to cluster $m$. The impact on the objective when assigning $i$ to $m$ is $M_{i,m}-M_{i,p}$, where $p$ is the other cluster. Since this difference is always positive when $i$ is placed in its most similar cluster, assigning $i$ as non-neutral always improves the objective. Then, since all objects are non-neutral, the problem is equivalent to the $k$-CC problem where we know max agreement and the full objective are equivalent (from Prop. 2). For $k>2$, this reasoning no longer holds, as contributions from other clusters can outweigh the within-cluster similarity to the most similar cluster (i.e., making the total contribution negative), potentially making neutrality optimal. However, we note that in our final objective (Eq. 3), when $\alpha$ and $\beta$ are involved, all terms will contribute with unique information even for $k=2$.
@@ Appendix C Impact of $\alpha$ and $\beta$
By examining the gradient in Eq. 8, we observe that the dominant term involving $\beta$ is $-\beta|S_{m}|$. Consequently, making $\beta$ large and negative increases the incentive to assign objects to non-neutral clusters. Moreover, since $-\beta|S_{m}|$ scales with cluster size, the local search procedure will favor placing an object $i$ in the largest non-neutral cluster. If $\beta$ is sufficiently large and negative, this term will completely dominate the objective, ensuring that no object is assigned to the neutral set (as the contribution to all non-neutral clusters remains positive). Ultimately, all objects will be placed in the largest non-neutral cluster.
Similarly, if $\beta$ is made very large and positive, $-\beta|S_{m}|$ will eventually dominate the objective, making the contribution to every non-neutral cluster negative for all objects. As a result, all objects will be assigned to the neutral set. ∎
From Prop. 4, we understand the extreme cases of $\beta$: (a) a small negative $\beta$ results in a maximally imbalanced non-neutral clustering (i.e., all objects in one non-neutral cluster), while (b) a large positive $\beta$ makes all objects neutral. For intermediate $\beta\in[\xi_{1},\xi_{2}]$, we analyze the gradient in Eq. 8. Increasing $\beta$ strictly reduces the contribution of object $i$ to each cluster $m\in[k]$, but since the term $-2\beta|S_{m}|$ scales with cluster size, larger clusters become less favorable, promoting balance. If $\beta$ is large enough, it forces $G_{i,m}<0$ for all $m\in[k]$, making neutrality optimal for object $i$. Note that this is more likely for low-degree objects, implying that high-degree objects (with clear cluster assignment) are more likely to remain non-neutral, resulting in dense non-neutral clusters. Consequently, increasing $\beta$ leads to smaller (i.e., more neutral objects) and denser non-neutral clusters, while maintaining balanced, as desired.
The parameter $\alpha$ has been studied in prior work [12, 44]. From Eq. 3, $\alpha$ balances maximizing intra-similarities and minimizing inter-similarities, which translates to a trade-off between cohesion within clusters and separation between them. A heuristic choice of $\alpha=1/(k-1)$ was proposed in [44], based on the observation that the number of intra-similarities scale linearly with $k$, while the number of inter-similarities grow quadratically. This choice prevents inter-similarities from dominating the objective. Finally, the term $-\alpha\phi_{i}$ indicates that $\alpha$ influences whether object $i$ becomes neutral, underscoring the need to account for inter-similarities in the objective (as suggested in Section 2.2).
@@ Appendix D Limitations of Polarity
[…]
@@ Appendix F Comparison with $\gamma$-polarity Objective from [21]
In contrast, our objective promotes balanced clustering through an additive regularization term which supports an arbitrary number of clusters $k$. A key advantage of additive regularization is its simplicity: it effectively corresponds to shifting intra-cluster similarities by $-\beta$ (see Prop. 1). This straightforward modification enables us to establish a linear convergence rate (Thm. 3) and significantly improves computational efficiency (see Section 3.4). In comparison, obtaining similar theoretical guarantees and efficiency via local search on polarity-based objectives, including $\gamma$-polarity, is likely infeasible or substantially more complex.
@@ G.4 Results on Synthetic Datasets with Imbalanced Clusters
The results are presented in Fig. 2. The noise level is fixed to $\eta=0.4$. We observe that our method remains robust as the cluster size imbalance increases, while the performance of baseline methods deteriorates more rapidly. This suggests that our approach does not rely on strict balance assumptions and is capable of effectively recovering imbalanced ground-truth clusterings.
G.5 Runtime Comparison of Local Search Improvements
Figure 3: Comparison of runtime for the three implementations of Alg. 2 (local search) introduced in Section 3.4 by varying the graph size $n$ and the number of non-neutral clusters $k$, using data generated from the m-SSBM model. The noise level is fixed at $\eta=0.4$. When varying $n$, we fix $k=4$; when varying $k$, we fix $n=5000$.
In Section 3.4, we introduced three different implementations of the local search procedure in Alg. 2: (i) the naive approach, referred to as LSNaive, which has a computational complexity of $O(k^{2}n^{2})$; (ii) an improved version that leverages the gradient, reducing the complexity to $O(Tkn)$, which we call LSImproved; and (iii) the final optimized version—used as LSPCD in all our experiments—presented in Alg. 3, which achieves a complexity of $O(kn^{2}+T(n+k))$.
Fig. 3 presents the runtime comparison as we vary the number of nodes $n$ and the number of non-neutral clusters $k$, using datasets generated from the m-SSBM model (Section G.3) with fixed noise level $\eta=0.4$. We fix $k=4$ when varying $n$, and $n=5000$ when varying $k$. LSNaive is included only for small values of $n$ due to its high computational cost, as shown in the leftmost plot. Across the remaining plots, the method in Algorithm 3 (LSPCD) demonstrates superior scalability with respect to both $n$ and $k$. While these are moderate-sized graphs, the performance gap would widen significantly for larger instances.
G.6 Aspects to Assess Solution Quality
@@
[…]
    Its range is $[-1,1]$, where $-1$ indicates that all non-zero intra-cluster similarities are negative, and $+1$ indicates that all are positive.
@@
[…]
    Its range is $[-1,1]$, where $-1$ indicates that all non-zero inter-cluster similarities are positive, and $+1$ indicates that all are negative.
@@
G.7 Varying $\alpha$ and $\beta$
Fig. 4 shows the effect of varying $\beta$. Very small or large $\beta$ values lead to poorer polarity, as they produce clustering solutions with too many or too few non-neutral objects, respectively. In contrast, intermediate $\beta$ values consistently yields competitive polarity with the best performing methods, while being more balanced.
In Fig. 5, we illustrate the impact of varying $\alpha$ and $\beta$. According to Fig. 5(a), increasing $\alpha$ naturally balances intra-cluster cohesion and inter-cluster opposition. The size proportion, defined as the fraction of non-neutral objects in $V$, remains constant as $\alpha$ varies. Fig. 5(b) shows that increasing $\beta$ monotonically reduces the number of non-neutral objects, leading to denser clusters, as indicated by improved MAC and MAO scores. Notably, balance remains stable across different $\beta$ values, unlike the baseline SCG-MA.
@@ G.7 Varying $\alpha$ and $\beta$
Figure 5: Investigation of the impact of $\alpha$ and $\beta$.
G.8 Results on Real-World Datasets
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 586 0.176 223.406 2 0.261 0.098 0.757 0.509 0.511 0.769 0.249
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 155 0.648 29.022 2 0.2 0.143 0.94 0.969 0.199 0.211 2.203
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 1278 0.428 62.322 2 0.038 0.015 0.831 0.673 0.06 0.548 2.457
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 915 0.71 146.109 2 0.279 0.014 1.0 0.114 0.17 0.353 3.376
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 235 0.251 75.903 2 0.207 0.055 0.969 0.836 0.337 0.167 29.957
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 1876 0.825 190.8 2 0.055 0.128 0.871 0.997 0.108 0.242 72.368
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 2188 0.73 127.784 2 0.12 0.013 0.907 0.74 0.066 0.351 59.119
@@
$k$ SIZE IF POL K MAC MAO CC+ CC- DENS ISO TIME (s)
$2$ LSPCD (avg) 599 0.3 81.985 2 0.093 0.034 0.917 0.87 0.15 0.109 57.022

### v3

@@ 7 $N_{\text{intra}}^{+}-N_{\text{inter}}^{+}$ 2 0 -2 0 -2
8 $N_{\text{intra}}^{-}+N_{\text{inter}}^{-} -$1 1 -1 1 1
9 $N_{\text{intra}}^{-}-N_{\text{inter}}^{+}$ ($f^{\text{MinDisagree}}$) -1 -1 -3 -1 -2
10 $N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ ($f^{\text{MinCut}}$) 0 0 -2 0 -1
@@ 13 $N_{\text{intra}}^{+}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$ 2 1 -2 1 -1
14 $N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+} -$1 0 -3 0 -1
@@
Now, consider a clustering with $k$ non-neutral clusters, where all intra-cluster similarities are +1+1 and all inter-cluster similarities are 1-1. If an unassigned object $i\in V$ has similarity +1+1 to all others, the max correlation objective (Eq. 13) assigns it to the largest non-neutral cluster, while the minimum cut objective (Eq. 14) keeps it neutral. For $k>2$, the full objective (Eq. 1) places $i$ in the largest non-neutral cluster if its size exceeds the sum of all others; otherwise, it remains neutral. Conversely, if object $i$ has similarity 1-1 to all others, max correlation keeps it neutral, whereas minimum cut assigns it to the smallest non-neutral cluster. For $k>2$, the full objective may assign $i$ as neutral or non-neutral depending on cluster sizes.
@@
For $k=2$, the full objective always increases (or remains constant) when an object is made non-neutral, aligning it with max agreement. To see this, consider a clustering with $k=2$ non-neutral clusters, and let $M_{i,m}=\sum_{j\in S_{m}}A_{i,j}$ be the total similarity of object $i$ to cluster $m$. The impact on the objective when assigning $i$ to $m$ is $M_{i,m}-M_{i,p}$, where $p$ is the other cluster. Since this difference is always positive when $i$ is placed in its most similar cluster, assigning $i$ as non-neutral always improves the objective. Then, since all objects are non-neutral, the problem is equivalent to the $k$-$C$ problem where we know max agreement and the full objective are equivalent (from Prop. 2). For $k>2$, this reasoning no longer holds, as contributions from other clusters can outweigh the within-cluster similarity to the most similar cluster (i.e., making the total contribution negative), potentially making neutrality optimal. However, we note that in our final objective (Eq. 3), when $\alpha$ and $\beta$ are involved, all terms will contribute with unique information even for $k=2$.
@@ Appendix C Impact of $\alpha$ and $\beta$
By examining the gradient in Eq. 9, we observe that the dominant term involving $\beta$ is $\beta|S_{m}|$. Consequently, making $\beta$ large and negative increases the incentive to assign objects to non-neutral clusters. Moreover, since $\beta|S_{m}|$ scales with cluster size, the local search procedure will favor placing an object $i$ in the largest non-neutral cluster. If $\beta$ is sufficiently large and negative, this term will completely dominate the objective, ensuring that no object is assigned to the neutral set (as the contribution to all non-neutral clusters remains positive). Ultimately, all objects will be placed in the largest non-neutral cluster.
Similarly, if $\beta$ is made very large and positive, $\beta|S_{m}|$ will eventually dominate the objective, making the contribution to every non-neutral cluster negative for all objects. As a result, all objects will be assigned to the neutral set. ∎
From Prop. 4, we understand the extreme cases of $\beta$: (a) a small negative $\beta$ results in a maximally imbalanced non-neutral clustering (i.e., all objects in one non-neutral cluster), while (b) a large positive $\beta$ makes all objects neutral. For intermediate $\beta\in[\xi_{1},\xi_{2}]$, we analyze the gradient in Eq. 9. Increasing $\beta$ strictly reduces the contribution of object $i$ to each cluster $m\in[k]$, but since the term $2\beta|S_{m}|$ scales with cluster size, larger clusters become less favorable, promoting balance. If $\beta$ is large enough, it forces $G_{i,m}<0$ for all $m\in[k]$, making neutrality optimal for object $i$. Note that this is more likely for low-degree objects, implying that high-degree objects (with clear cluster assignment) are more likely to remain non-neutral, resulting in dense non-neutral clusters. Consequently, increasing $\beta$ leads to smaller (i.e., more neutral objects) and denser non-neutral clusters, while maintaining balanced, as desired.
The parameter $\alpha$ has been studied in prior work [12, 44]. From Eq. 3, $\alpha$ balances maximizing intra-similarities and minimizing inter-similarities, which translates to a trade-off between cohesion within clusters and separation between them. A heuristic choice of $\alpha=1/(k-1)$ was proposed in [44], based on the observation that the number of intra-similarities scale linearly with $k$, while the number of inter-similarities grow quadratically. This choice prevents inter-similarities from dominating the objective. Finally, the term $\alpha\phi_{i}$ indicates that $\alpha$ influences whether object $i$ becomes neutral, underscoring the need to account for inter-similarities in the objective (as suggested in Section 2.2).
@@ Appendix D Limitations of Polarity
To illustrate the limitation of polarity (Eq 2), we refer to Example 2 from [21], which considers a signed graph with 12 objects: $\{\texttt{A},\texttt{B},\texttt{C},\texttt{D},\texttt{E},\texttt{F},\texttt{G},\texttt{H},\texttt{I},\texttt{J},\texttt{K},\texttt{L}\}$. The sign of each similarity can be found in Figure 3 of [21]. The study evaluates the following three clustering solutions:
  • $S^{(1)}=\{\{\texttt{A},\texttt{B},\texttt{C},\texttt{D}\},\{\texttt{E},\texttt{F},\texttt{G},\texttt{H}\}\}$
  • $S^{(2)}=\{\{\texttt{A},\texttt{B},\texttt{C},\texttt{D}\},\{\texttt{E},\texttt{F},\texttt{G},\texttt{H},\texttt{I},\texttt{J},\texttt{K},\texttt{L}\}\}$
  • $S^{(3)}=\{\emptyset,\{\texttt{E},\texttt{F},\texttt{G},\texttt{H},\texttt{I},\texttt{J},\texttt{K},\texttt{L}\}\}$
@@ Appendix F Comparison with $\gamma$-polarity Objective from [21]
In contrast, our objective promotes balanced clustering through an additive regularization term which supports an arbitrary number of clusters $k$. A key advantage of additive regularization is its simplicity: it effectively corresponds to shifting intra-cluster similarities by $\beta$ (see Prop. 1). This straightforward modification enables us to establish a linear convergence rate (Thm. 3) and significantly improves computational efficiency (see Section 3.4). In comparison, obtaining similar theoretical guarantees and efficiency via local search on polarity-based objectives, including $\gamma$-polarity, is likely infeasible or substantially more complex.
@@ G.4 Results on Synthetic Datasets with Imbalanced Clusters
The results are presented in Figure 3. The noise level is fixed to $\eta=0.4$. We observe that our method remains robust as the cluster size imbalance increases, while the performance of baseline methods deteriorates more rapidly. This suggests that our approach does not rely on strict balance assumptions and is capable of effectively recovering imbalanced ground-truth clusterings.
G.5 Runtime Comparison on Large Scale Synthetic Data
Figure 4: Runtime comparison on large-scale synthetic datasets generated using the m-SSBM model. LSPCD consistently achieves higher F1-scores than SCG-MA while requiring less runtime, demonstrating superior scalability and efficiency. We fixed $\eta=0.45$, $k=6$, and $\ell=500$.
To further highlight the scalability of our approach, we present additional experiments on large synthetic datasets generated using the m-SSBM model described in the paper, containing up to 250,000 objects and 150 million edges with non-zero edge weight. Our method is capable of handling even larger datasets; the only constraint was the memory required to store the matrix of pairwise relations on the machine used for our experiments, a limitation that could be easily overcome by using machines with larger memory capacities. We compare LSPCD with the strongest baseline, SCG with min-angle rounding (SCG-MA), and report the F1-score (ground-truth cluster recovery) and runtime in seconds. Other baselines yield lower F1-scores, and most are also slower in runtime. We fix $\eta=0.45$, $k=6$ and $\ell=500$.
The results are shown in Figure 4. Our method consistently outperforms SCG in terms of F1 score while also being more efficient, demonstrating its ability to scale to larger graphs.
G.6 Impact of $\xi$ on Imbalance Factor
Evaluating unsupervised learning methods in the absence of ground truth is inherently challenging, and there is often no universally or uniquely accepted metric or criterion. As a result, it is common practice to report multiple metrics, each reflecting a different objective or perspective. In this study, for real-world datasets, we report both polarity and imbalance factor. As shown in our extensive experiments, our method is the only one that consistently performs well w.r.t. both metrics across various datasets.
Given the difficulty of evaluation in such settings, no single fixed value of $\xi$ can be considered optimal. In the absence of prior preference, a reasonable approach is to examine multiple values of $\xi$ and analyze how different methods perform under each. We experimented with a range of values and found that the conclusions remain stable. As mentioned in the main paper, we chose $\xi=3$ since it provides a sharper distinction between solutions with different degrees of balance, compared to $\xi=1$, which would have been the natural choice otherwise as it corresponds to Shannon entropy. To further demonstrate robustness, we now also report results for $\xi=1,3,4$ for a subset of methods (see Table 3). Across all tested values, the relative ranking of methods remains consistent, and our conclusions are unchanged (again when SCG is more balanced, it usually yields low polarity). Results are consistent for even larger values of $\xi$.
Table 3: The imbalance factor (IF) for three different values of $\xi=1,3,4$.
REF SD WikiC EP WikiP
$k$ Method $\xi{=}1$ $\xi{=}3$ $\xi{=}4$ $\xi{=}1$ $\xi{=}3$ $\xi{=}4$ $\xi{=}1$ $\xi{=}3$ $\xi{=}4$ $\xi{=}1$ $\xi{=}3$ $\xi{=}4$ $\xi{=}1$ $\xi{=}3$ $\xi{=}4$
2 LSPCD (ours) 0.88 0.78 0.66 0.50 0.31 0.22 0.93 0.87 0.79 0.89 0.80 0.68 0.56 0.37 0.27
SCG-MA 0.05 0.02 0.01 0.06 0.02 0.01 0.77 0.62 0.48 0.13 0.05 0.04 0.06 0.02 0.01
SCG-MO 0.05 0.02 0.01 0.04 0.01 0.01 0.69 0.51 0.39 0.13 0.05 0.03 0.03 0.01 0.01
SCG-B 0.11 0.04 0.03 0.15 0.06 0.04 0.84 0.72 0.59 0.14 0.06 0.04 0.13 0.05 0.04
4 LSPCD (ours) 0.58 0.46 0.38 0.74 0.61 0.50 0.81 0.65 0.50 0.78 0.65 0.54 0.55 0.34 0.24
SCG-MA 0.75 0.71 0.66 0.51 0.31 0.22 0.18 0.08 0.05 0.50 0.36 0.27 0.64 0.55 0.49
SCG-MO 0.75 0.72 0.68 0.41 0.25 0.17 0.50 0.30 0.21 0.54 0.40 0.30 0.46 0.35 0.28
SCG-B 0.00 0.00 0.00 0.55 0.43 0.35 0.95 0.90 0.82 0.68 0.60 0.51 0.34 0.25 0.19
6 LSPCD (ours) 0.49 0.37 0.31 0.66 0.50 0.40 0.80 0.62 0.48 0.71 0.55 0.43 0.76 0.61 0.49
SCG-MA 0.71 0.66 0.60 0.35 0.21 0.15 0.22 0.09 0.06 0.63 0.55 0.50 0.53 0.45 0.40
SCG-MO 0.78 0.74 0.71 0.56 0.37 0.26 0.50 0.28 0.19 0.42 0.31 0.24 0.36 0.27 0.22
SCG-B 0.00 0.00 0.00 0.45 0.36 0.29 0.83 0.76 0.67 0.53 0.46 0.39 0.26 0.19 0.14
G.7 Aspects to Assess Solution Quality
@@
    $\texttt{CC+}=\frac{N^{+}_{\text{intra}}-N^{-}_{\text{intra}}}{N^{+}_{\text{intra}}+N^{-}_{\text{intra}}}.$
    Its range is [1,1][-1,1], where 1-1 indicates that all non-zero intra-cluster similarities are negative, and +1+1 indicates that all are positive.
@@
    $\texttt{CC-}=\frac{N^{-}_{\text{inter}}-N^{+}_{\text{inter}}}{N^{-}_{\text{inter}}+N^{+}_{\text{inter}}}$.
    Its range is [1,1][-1,1], where 1-1 indicates that all non-zero inter-cluster similarities are positive, and +1+1 indicates that all are negative.
@@
G.8 Varying $\alpha$ and $\beta$
Figure 5 shows the effect of varying $\beta$. Very small or large $\beta$ values lead to poorer polarity, as they produce clustering solutions with too many or too few non-neutral objects, respectively. In contrast, intermediate $\beta$ values consistently yields competitive polarity with the best performing methods, while being more balanced.
In Figure 6, we illustrate the impact of varying $\alpha$ and $\beta$. According to Figure 6(a), increasing $\alpha$ naturally balances intra-cluster cohesion and inter-cluster opposition. The size proportion, defined as the fraction of non-neutral objects in $V$, remains constant as $\alpha$ varies. Figure 6(b) shows that increasing $\beta$ monotonically reduces the number of non-neutral objects, leading to denser clusters, as indicated by improved MAC and MAO scores. Notably, balance remains stable across different $\beta$ values, unlike the baseline SCG-MA.
@@ G.8 Varying $\alpha$ and $\beta$
Figure 6: Investigation of the impact of $\alpha$ and $\beta$ for LSPCD (Alg. 3).
G.9 Results on Real-World Datasets
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 586 0.176 223.406 2 0.261 0.098 0.757 0.509 0.511 0.769 0.249
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 155 0.648 29.022 2 0.2 0.143 0.94 0.969 0.199 0.211 2.203
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 1278 0.428 62.322 2 0.038 0.015 0.831 0.673 0.06 0.548 2.457
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 915 0.71 146.109 2 0.279 0.014 1.0 0.114 0.17 0.353 3.376
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 235 0.251 75.903 2 0.207 0.055 0.969 0.836 0.337 0.167 29.957
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 1876 0.825 190.8 2 0.055 0.128 0.871 0.997 0.108 0.242 72.368
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 2188 0.73 127.784 2 0.12 0.013 0.907 0.74 0.066 0.351 59.119
@@
$k$ SIZE IF POL K MAC MAO CC+ $C$- DENS ISO TIME (s)
22 LSPCD (avg) 599 0.3 81.985 2 0.093 0.034 0.917 0.87 0.15 0.109 57.022